OpenStudy (anonymous):

CAN ANYONE HELP ME??????? b=-1 (x,y)=(-3,0) f(x)=mx+b I have to write a rule for the linear equation .

OpenStudy (anonymous):

So if you graph the b and the point you were given, you can find the slope. I did this and got a slope of -1/3. Then I substituted the variables. So you now have f (-3) = -1/3 (-3) -1 Now you solve and come to get f (-3) = 0 I'm not sure if this is what you are looking for, but this is what I did.

OpenStudy (anonymous):

Thanks but I am trying to find the rule of the linear equation . My teacher gave me the answer which is f9x)= -1/3x-1 which is what u got but I have to show how I got that >_<

OpenStudy (anonymous):

Okay well here is my work f (-3) = -1/3 (-3) - 1 f (-3) = 1 - 1 f (-3) = 0

OpenStudy (anonymous):

Thanks I have another one that im stuck on could u help me possibly?

OpenStudy (anonymous):

Yeah i might be able to help. Let me see it

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay and what do you have to do?

OpenStudy (anonymous):

I have the answer which is the rule . it is f(x)=-2x +10 but I have to show the work on how to get that. :(

OpenStudy (anonymous):

Okay give me a few minutes and I will see if i can get it

OpenStudy (anonymous):

I tried it but didnt get the right answer, and I'm not sure how to do it exactly. Sorry :-\

OpenStudy (anonymous):

Its fine but thank u soooo much for what u have done ! <3 :)